Lindsay​ Electronics, a small manufacturer of electronic research​ equipment, has approximately 7,000items in its inventory and has hired Joan​ Blasco-Paul to manage its inventory. Joan has determined that 8​%of the items in inventory are A​ items, 38% are B​ items, and 54​%are C items. She would like to set up a system in which all A items are counted monthly​ (every 22 working​ days), all B items are counted quarterly​ (every 60 working​ days), and all C items are counted semiannually​ (every 119 working​ days). How many items need to be counted each​ day? The total number of items that need to be counted each day is ______items ​(round your response to the nearest whole​ number)
Answer #1

From the given data:-

A items are 0.08 * 7000 = 560

B items are 0.38 * 7000 = 2,660

C Items are 0.54 * 7000 = 3,780

Units to be counted everyday = 560/22 + 2,660/60 + 3,780/119 = 25.45 + 44.33 + 31.76 = 101.54 or 102

A thumbs up, pls !!